### v1

# A General Model for Static Contact Angles

Abstract
The problem of contact angle and hysteresis determination has direct implications for engineering applications of wetting, colloid and surface science. Significant technical challenges can arise under real-world operating conditions, because the static contact angle is strongly influenced by contamination at the liquid-solid and liquid-vapor interfaces, chemical aging over long times, and environmental variables such as relative humidity and temperature. Analytical models that account for these real-world effects are therefore highly desirable to guide the rational design of robust applications. This work proposes a unified and simple-to-use model that expands Young’s local thermodynamic approach to consider surfaces with topographic features of general geometry and varying degrees of liquid infiltration. The unified model recovers classical wetting limits (Wenzel, Cassie-Baxter, and hemiwicking), accounts for observable intermediate states (e.g., impregnating Cassie), and identifies a new limiting state with potential realizability: a Cassie state accompanied by a precursor film, termed the Inverse Wenzel state.
Keywords: Wetting, Static Contact Angle, Wetting Hysteresis, Wenzel State, Cassie-Baxter State, Hemiwicking.
1 Introduction
The determination of static contact angles on technologically relevant surfaces that exhibit micro- and nanoscale physical and chemical features remains a foundational problem in colloid and surface science.1; 2; 3; 4; 5 The problem of contact angle and hysteresis determination has direct implications for engineering applications ranging from surface wetting control,6; 7; 8; 9; 10; 11 capillary infiltration in porous materials,12; 13; 14; 15; 16; 17; 18 and particle adhesion at interfaces,19; 20; 21; 22; 23; 24 to additive manufacturing.25; 26; 27 The functionality and performance of devices across these applications are highly sensitive to how closely realized contact angles and hysteresis ranges agree with experimental and analytical predictions. Significant technical challenges can arise under real-world operating conditions, because the static contact angle, is strongly influenced by surface cleanliness and contamination at the liquid-solid and liquid-vapor interfaces, by chemical aging processes over long times, and by environmental variables such as relative humidity and temperature. Analytical models that account for these real-world effects are therefore highly desirable to guide the rational design of robust applications.
For an ideal surface that is perfectly flat and chemically homogeneous, and a volatile liquid in thermodynamic equilibrium with its vapor phase, the equilibrium contact angle is unique and given by Young’s law,28
$\gamma_{\ell v}\cos\theta_{Y}=\gamma_{sv}-\gamma_{s\ell},$ (1)
which relates the equilibrium (Young) contact angle $\theta_{Y}$ to the interfacial surface energies $\gamma_{s\ell}$, $\gamma_{\ell v}$, and $\gamma_{sv}$ of the solid-liquid, liquid-vapor, and solid-vapor interfaces, respectively. The interfacial surface energies in Eq. 1 are conventionally treated as material properties uniquely determined by the particular surface and fluid chemistry. Although simple tangential force-balance derivations of Eq. 1 for a plane surface have known flaws, rigorous thermodynamic arguments have established the validity of Young’s law for general geometric configurations.29; 30 Moreover, recent studies indicate that Young’s law remains applicable at unexpectedly small length scales, such as contact-line curvature and surface feature dimensions, down to roughly ten liquid molecule diameters.31; 32; 33
Engineered or natural surfaces with micro/nanoscale roughness or physical topography, exhibit a range of static contact angles that can significantly differ from the Young angle (Eq. 1) and are observed as the contact line quasi-statically recedes or advances. To account for the observed static contact angles and hysteresis phenomenon, two wetting models have been widely adopted: the Wenzel model34
$\cos\theta_{W}=r\,\cos\theta_{Y},$ (2)
and the Cassie-Baxter model35
$\cos\theta_{CB}=\varphi_{S}\,\cos\theta_{Y}-\big(1-\varphi_{S}\big),$ (3)
for the apparent equilibrium contact angles $\theta_{W}$ and $\theta_{CB}$ corresponding to each state, $r\geq 1$ is ratio of the actual wetted solid area to its projected area, and $\varphi_{S}$ is the planar solid area fraction in contact with the wetting liquid phase. It is important to note that the Wenzel and Cassie-Baxter models in Eqs. 23 represent two limiting wetting states: full liquid impregnation of the surface topography (Wenzel state) and strictly zero liquid infiltration beneath the liquid phase (Cassie-Baxter state). Notably, intermediate wetting states with partial and localized infiltration of the surface topography are realized under typical ambient conditions, and these commonly correspond to metastable configurations separated by large but finite free-energy barriers.36; 37; 38; 39; 40; 41 In particular, certain combinations of micro/nanoscale topography and low intrinsic Young angle, promote stable wetting states with the infiltration or liquid ahead of the contact line, through the phenomenon of hemiwicking.42; 43; 44; 45
Prior work has focused on developing unified models that bridge the pure Wenzel and Cassie-Baxter extremes and extend these classical models.46; 47; 48 These efforts employ free energy minimization to account for multiple wetting states and assess their thermodynamic stability, for specific wetting configurations and surface geometries such as a macroscale sessile droplets on much smaller micropillar arrays. Despite the significant progress on the fundamental and applied aspects, open questions remain, namely how to predict static contact angles and the hysteresis range from a knowledge of topographic parameters, surface chemistry aging, and ambient conditions affecting the degree of liquid infiltration on the surface below and ahead of the wetting liquid phase. In an effort to address this, this work proposes a general and simple-to-use model, expanding Young’s local thermodynamic approach to consider surfaces with topographic features of general geometry and varying degrees of liquid infiltration. The introduced unified model recovers classical wetting limits (Wenzel, Cassie-Baxter, and hemiwicking states), accounts for observable intermediate states (e.g., impregnating Cassie), and further identifies a new limiting state with potential realizability: a Cassie state accompanied by a precursor film, here termed the Inverse Wenzel state.
2 Theoretical Derivation
The derivation presented in this section is valid for predicting the local static contact angle when a volatile liquid wetting a solid surface is in chemical, thermal, and mechanical equilibrium with an ambient air/vapor phase. As illustrated in Fig. 1a, the liquid phase lies on the Bulk (B) side of the contact line and the vapor phase on the Ambient (A) side. We consider a chemically homogeneous substrate with (random or regular) three-dimensional micro- and nanoscale topographic features (Fig. 1a) producing a full surface area $S\geq A$, larger than the projected surface area $A$.
Geometric parameters.
The full surface area is generally decomposed as $S=S_{\text{top}}+S_{\text{lat+bot}}$, where $S_{\text{top}}\!\geq 0$ is the planar area of feature tops, if present at all, and $S_{\text{lat+bot}}\!>0$ denotes the lateral-plus-basal area generated by the topography. We thus define the solid-top area fraction $\varphi_{S}=S_{\text{top}}/A$ and the lateral-basal area fraction $\Lambda=S_{\text{lat+bot}}/A$. With this notation, the Wenzel roughness ratio in Eq. 2 is $r=S/A=\varphi_{S}+\Lambda$. We will define $\varphi_{L}(h)=A_{\ell v}/A$ as the area fraction that would correspond to liquid-vapor interface if the surface topography were fully infiltrated by liquid up to the reference height $h$ of micro- or nanoscale dimensions. We will additionally assume that the reference liquid height $h$ within the infiltrated topography sets the local contact line position and is nearly constant within small distances $\lambda\sim h$ from the contact line. Note that, under assumptions leading to the Cassie-Baxter model (Eq. 3), we have $\varphi_{L}=1-\varphi_{S}$ but this equality is strictly valid when all topographic features have approximately the same height $h_{r}$ and the liquid-vapor interface is perfectly flat and coincides with the reference level $h=h_{r}$; see Ref. 49 for a detailed discussion. For random self-affine micro/nanoscale topography with tall peaks of small lateral extent protruding above the r.m.s. height, one may approximate $\varphi_{L}\simeq 1-\varphi_{S}$; an error on the order of roughly $15\%$ can be estimated for Gaussian height statistics.
Figure 1: Generalized model for static contact angles on a surface with micro/nanoscale surface topography. (a) General wetting state and unified model for varying degrees of liquid infiltration in the bulk side ($0\leq\phi_{B}\leq 1$) and ambient side ($0\leq\phi_{A}\leq 1$) of the contact line. (b) Wenzel state corresponding to full liquid infiltration under the liquid phase ($\phi_{B}=0$) and a perfectly dry surface under the ambient phase ($\phi_{A}=1$). (c) Cassie-Baxter state corresponding to strictly no liquid infiltration under the liquid and ambient phases ($\phi_{B}=\phi_{A}=1$). (d) Hemiwicking state corresponding to fill liquid infiltration under the liquid and ambient phases ($\phi_{B}=\phi_{A}=0$). (e) Inverse Wenzel state corresponding to strictly no liquid infiltration under the liquid phase ($\phi_{B}=1$) and full liquid infiltration under the ambient phase ($\phi_{A}=0$).
Generalized wetting conditions.
Departing from the conventional treatment that leads to the classical Wenzel and Cassie-Baxter states, we consider that in general wetting and infiltration of the surface topography may not be ideally uniform; near the contact line, localized air/vapor voids may exist beneath the liquid phase, while localized liquid infiltration by wicking and condensation may occur through the texture on the ambient side exposed to the ambient vapor phase (see Fig. 1a). Hence, let $\phi_{B}\in[0,1]$ denote the averaged fraction of surface features not accessed by the liquid or partially occupied by air/vapor in the bulk side of the contact line and $\phi_{A}\in[0,1]$ denote the averaged fraction of surface features accessed and partially filled by liquid in the ambient side of the contact line. The limit condition with zero liquid-vapor interfacial area $S_{\ell v}=0$ on the bulk and ambient side thus correspond to $\phi_{R}=0$ and $\phi_{A}=1$; i.e., the Wenzel state (Fig. 1b). Similarly, the limit with $\phi_{R}=1$ and $\phi_{A}=1$ corresponds to the case that the vapor-liquid interface is $S_{\ell v}=\varphi_{L}A$; i.e., the Cassie-Baxter state (Fig. 1c).
Free-energy change and stationarity.
Following a conventional thermodynamic argument, we will consider that receding or advancing displacements of the contact line, under constant temperature, volume, and chemical equilibrium, produce a change in the free energy (or grand potential)42; 50 $dF=\gamma_{\ell v}\cos\theta+\gamma_{s\ell}dS_{s\ell}+\gamma_{sv}dS_{sv}+\gamma_{\ell v}dS_{\ell v}$, where $\gamma_{s\ell}$, $\gamma_{sv}$, and $\gamma_{\ell v}$ are the solid-liquid, solid-vapor, and liquid-vapor interfacial surface energies, and $S_{s\ell}$, $S_{sv}$, and $S_{\ell v}$ are the corresponding interfacial surface areas. Based on the geometric considerations above, a local displacement of the contact line over a projected area differential $\Delta A$, produces the following interfacial areas change:
$\Delta S_{s\ell}$ $=\Big(\,\varphi_{S}\;+\;\Lambda(\phi_{A}-\phi_{B})\,\Big)\Delta A,$ (4)
$\Delta S_{sv}$ $=\Big(\,-\varphi_{S}\;+\;\Lambda(\phi_{A}-\phi_{B})\,\Big)\Delta A,$ (5)
$\Delta S_{\ell v}$ $=-\,\Delta A\,\varphi_{L}\,(1-\phi_{A}-\phi_{B}).$ (6)
By imposing free-energy stationarity $dF=0$, invoking the Young contact angle definition in Eq. 1, and introducing the surface area changes in Eqs. 45, one obtains that the contact angle for static equilibrium must satisfy
$\cos\theta\;=\;\cos\theta_{Y}\,\Big(\varphi_{S}+\Lambda(\phi_{A}-\phi_{B})\Big)\;+\;\varphi_{L}\,\big(1-\phi_{A}-\phi_{B}\big).\;$ (7)
The general model for the local contact angle derived in Eq. 7 recovers the limiting behaviors predicted by previous well-established models, by making the conventional assumption that $\varphi_{L}+\varphi_{S}=1$: in the Wenzel state with $\phi_{B}=0$ and $\phi_{A}=1$, Eq. (7) reduces to Eq. 2 (Fig. 1b); in the Cassie-Baxter state for which $\phi_{B}=1$ and $\phi_{A}=1$ and the surface topography is fully filled by air/vapor on both the bulk and ambient sides, Eq. 7 yields Eq. 3 (Fig. 1c); and in the hemiwicking configuration (Fig. 1d), for which the surface topography is fully infiltrated by liquid on both sides adjacent to the contact line ($\phi_{B}=0$ and $\phi_{A}=0$), Eq. (7) simplifies to $\cos\theta=\varphi_{S}\cos\theta_{Y}+(1-\varphi_{S})$.42 In addition, Eq. 7 predicts a fourth limiting state, here termed the inverse Wenzel state (Fig. 1e), for which $\phi_{B}=1$ and $\phi_{A}=0$ and Eq. 7 gives
$\cos\theta\;=\;\cos\theta_{Y}(\varphi_{S}-\Lambda).$ (8)
This inverted state, which can, for example, yield large local contact angles $\theta\gtrsim 90^{\circ}$ on a hydrophilic substrate with $\theta_{Y}<90^{\circ}$, can be realized when the surface topography traps a localized metastable vapor void next to a region highly infiltrated by liquid.
3 Results and Discussion
This section presents analytical results based on Eq. 7 for substrates with varying wettability, as characterized by the Young angle $\theta_{Y}$, and different surface topographies produced by random and engineered features. The analysis covers the full range of surface wetting conditions characterized by the liquid infiltration fractions on both the bulk and ambient sides adjacent to the contact line. For simplicity, the results reported here (Figs. 2–4) assume $\varphi_{L}=1-\varphi_{S}$. The implications of this approximation are discussed in the conclusions section.
Figure 2: Sharp random topography. Local static contact angle $\theta$  predicted by Eq. 7 as a function of liquid-vapor surface fractions $\phi_{B}$ (bulk) and $\phi_{A}$ (ambient) adjacent to the contact line, for three Young angles: $\theta_{Y}=40^{\circ},\,90^{\circ},\,\&\,120^{\circ}$. The results correspond to surface topography with a small solid-top area fraction $\varphi_{S}=0.01$ and for two different lateral-plus-basal area fractions $\Lambda=1$ (top row) and $\Lambda=2$ (bottom row). Wetting states: (W) Wenzel, (CB) Cassie-Baxter, (IW) Inverse Wenzel, and (H) Hemiwicked.
Sharp random topography.
The results in Fig. 2 correspond to surfaces with sharp topographic features, with a small solid-top area fraction $\varphi_{S}=0.01$ and moderate-to-large lateral-plus-basal area fractions $\Lambda=1$ and $\Lambda=2$. The combination of topographic parameters in Fig. 2 represents, for example, common glass surfaces with natural micro/nanoscale roughness16; 51; 18 or engineered nanostructures with random nanomaterial deposition.52; 53 The analytical results in Fig. 2 show that a wide range of static contact angles, which deviate largely from the intrinsic Young contact angle, is feasible for this type of surface topography. For sufficiently “tall” topographic features giving large lateral area fractions $\Lambda\geq 2$, robust superhydrophilicity with $\theta\to 0$ can be realized for $\phi_{B}=0$ and $0\leq\phi_{A}\leq 1$ on wettable substrates with $\theta\lesssim 40^{\circ}$. This indicates that superwetting with vanishing hysteresis can be attained regardless of the conditions on the surface exposed to the ambient phase. Notably, large lateral area fractions $\Lambda\geq 2$ in a hydrophilic substrate can promote superhydrophobic states with $\theta\geq 150^{\circ}$ if the liquid phase is nearly fully suspended under topographically trapped air ($\phi_{B}\geq 0.9$), for arbitrary wetting conditions on the ambient side ($0\leq\phi_{A}\leq 1$). This surprising result correspond to the so-called inverse Wenzel state identified in Fig. 1e.
Engineered pillared structures.
Figure 3: Engineered pillared structures. Local static contact angle $\theta$  predicted by Eq. 7 as a function of liquid-vapor surface fractions $\phi_{B}$ (bulk) and $\phi_{A}$ (ambient) adjacent to the contact line, for three Young angles: $\theta_{Y}=40^{\circ},\,90^{\circ},\,\&\,120^{\circ}$. The results correspond to surfaces with sharp topographic features, with a small solid-top area fraction $\varphi_{S}=0.5$ and for two different lateral-plus-basal area fractions $\Lambda=1$ (top row) and $\Lambda=2$ (bottom row). Wetting states: (W) Wenzel, (CB) Cassie-Baxter, (IW) Inverse Wenzel, and (H) Hemiwicked.
The static contact angle predictions in Fig. 3 correspond to surface topographies with a moderately large top-solid area fraction, $\varphi_{S}=0.5$, and lateral-plus-basal area fractions, $\Lambda=1$ and $\Lambda=2$. This type of surface topography is typically attained through engineered surface structures of micro- or nanoscale dimensions, featuring flat top surface areas with controlled dimensions and lateral-basal surface areas determined by the feature height and period (see, for example, Refs. 54; 9; 10). In this case, analytical predictions show less sensitivity to variations in the wetting conditions on the surface exposed to the liquid bulk and ambient vapor phase, resulting in relatively moderate hysteresis ranges for Young angles, $\theta_{Y}$, between 60 and 120, and shallow surface features for which $\Lambda\simeq 1$ (cf. Fig. 3). As shown in Fig. 3, surfaces with a moderately large top-solid area fraction, $\varphi_{S}=0.5$, require low Young angles, $\theta_{Y}\lesssim 40^{\circ}$, and tall structures with short periods that yield large lateral-basal area fractions, $\Lambda\gtrsim 2$, to produce robust superhydrophilicity for varying conditions on the surface side exposed to the ambient phase. The realizability of large contact angles, $\theta\geq 150^{\circ}$, for hydrophilic substrates with $\theta_{Y}\simeq 40^{\circ}$ is confined to a very narrow range of conditions ($\phi_{B}>0.9$ and $\phi_{A}<0.1$), for which the so-called inverse Wenzel state can exist. In addition, sticky superhydrophobicity with $\theta\geq 150^{\circ}$ can be attained on this type of surface topography (cf. Fig. 3) when the substrate is highly hydrophobic ($\theta_{Y}\geq 120^{\circ}$) and the liquid bulk infiltrates the topography underneath while the ambient side is partially wetted ($\phi_{B}\leq 0.5$ and $\phi_{A}\geq 0.9$).
Micro/nanocavity arrays.
The final cases analyzed (cf. Fig. 4) correspond to a large top-solid flat area fraction, $\varphi_{S}=0.8$, and lateral-plus-basal area fractions $\Lambda=1$ and $\Lambda=2$. This type of surface topography can be attained, for example, by patterning a smooth and flat surface with arrays of cavities with controllable micro- or nanoscale dimensions (e.g., Refs. 55; 56; 57). Note that if the cavities are not interconnected, capillary condensation and/or direct prewetting of the surface, rather than hemiwicking, are feasible mechanisms for infiltrating the surface topography. For this type of topography, robust superhydrophilicity with $\theta\to 0$ can be realized for $\phi_{A}\gtrsim 0.5$ and $\phi_{B}\lesssim 0.5$ in the case of large lateral-basal areas with $\Lambda>1$ and hydrophilic substrates with $\theta_{Y}\lesssim 40^{\circ}$. Superhydrophobicity is more difficult to achieve with this type of surface topography, and $\theta\geq 150^{\circ}$ is predicted only for highly hydrophobic substrates when $\phi_{A}\gtrsim 0.8$ and $\phi_{B}\lesssim 0.5$, which represent a rather narrow range of conditions.
Figure 4: Micro/nanocavity arrays. Local static contact angle $\theta$  predicted by Eq. 7 as a function of liquid-vapor surface fractions $\phi_{B}$ (bulk) and $\phi_{A}$ (ambient) adjacent to the contact line, for three Young angles: $\theta_{Y}=40^{\circ},\,90^{\circ},\,\&\,120^{\circ}$. The results correspond to surfaces with sharp topographic features, with a small solid-top area fraction $\varphi_{S}=0.8$ and for two different lateral-plus-basal area fractions $\Lambda=1$ (top row) and $\Lambda=2$ (bottom row). Wetting states: (W) Wenzel, (CB) Cassie-Baxter, (IW) Inverse Wenzel, and (H) Hemiwicked.
4 Conclusions
This work presents a unified and simply to use model for predicting static contact angles that extends Young’s local thermodynamic framework to surfaces with general topography and non-uniform local liquid infiltration. The formulation makes explicit the geometric controls through the solid-top area fraction $\varphi_{S}$ and the normalized lateral plus basal area $\Lambda$, and it resolves the local wetting state adjacent to the contact line with two side specific infiltration fractions $(\phi_{A},\phi_{B})$. In closed form, the model recovers the classical limits of Wenzel, Cassie-Baxter, and hemiwicking, and it predicts a previously unreported limiting configuration with practical relevance, an Inverse Wenzel state characterized by a Cassie core coexisting with a hemiwicked film. Because the inputs are measurable from surface geometry and standard force tensiometer data, the framework links morphology, local infiltration, and apparent angle, and it provides sharp bounds for advancing and receding states.
Analytical exploration across representative surface morphologies yields design guidelines. For sharp random relief with very small $\varphi_{S}$ and large $\Lambda$, the model predicts broad tunability: robust superhydrophilicity with $\theta\to 0$ on intrinsically wettable chemistries when the bulk side is liquid filled ($\phi_{B}=0$), with weak dependence on the ambient side condition. The same class of morphologies can also support very large apparent angles on hydrophilic substrates through the Inverse Wenzel state if the liquid is nearly fully suspended over trapped vapor on the bulk side ($\phi_{B}\gtrsim 0.9$), which highlights how large $\Lambda$ amplifies $\theta$. For engineered pillar arrays with moderate $\varphi_{S}\sim 0.5$, sensitivity to $(\phi_{A},\phi_{B})$ is reduced, so achieving superhydrophilicity requires taller features and low $\theta_{Y}$, while superhydrophobicity appears only in a narrow window of side specific infiltration. For cavity rich surfaces with large $\varphi_{S}\sim 0.8$, superhydrophilicity is attainable for hydrophilic chemistries when $\Lambda>1$ and the ambient side wets appreciably, whereas superhydrophobic states are limited to intrinsically hydrophobic chemistries and restricted ranges of $(\phi_{A},\phi_{B})$.
The model also strengthens connections to near equilibrium (quasi-static) dynamics. By predicting feasible ranges of apparent angles and hysteresis for given morphology, chemistry, and ambient conditions, it provides a basis to rationalize the slow quasi static spreading of droplets and capillary imbibition reported for glass and related materials 58; 16; 51; 59; 52; 53; 18. The side specific infiltration parameters $(\phi_{A},\phi_{B})$ capture metastable pinning and depinning pathways near the contact line, thus informing the control of entry pressures, relaxation times, and near-equilibrium flow thresholds. The same predictions constrain the capillary energy landscape that governs the quasi static relaxation of micro and nanoparticles at liquid fluid and liquid solid interfaces 20; 60; 61; 62; 63; 24; 64; 65, enabling surface design strategies, for example increasing $\Lambda$ at fixed $\varphi_{S}$, to accelerate or arrest adsorption and to tune dissipation during interfacial relaxation.
Future experimental works can use force versus displacement measurements with the Wilhelmy plate method to calibrate, test, and extend the model under controlled surface chemistry, cleanliness, ambient humidity, and temperature. The direct relation between the measured force and the apparent angle along advancing and receding branches enables inference of the side specific infiltration fractions $(\phi_{A},\phi_{B})$, validation of the predicted bounds and transitions, and targeted searches for the proposed Inverse Wenzel state. In parallel, the framework can be closed with thermodynamic relations that predict $\phi_{B}$ and $\phi_{A}$ from measurable quantities such as Laplace pressure, interfacial work, and local free energy differences, for example by minimizing the interfacial free energy of a representative unit cell at the prescribed reference height given $\theta_{Y}$, $\varphi_{S}$, and $\Lambda$. The combination of Wilhelmy data and these closures will enable parameter free predictions of static angles and hysteresis and guide refinement of the model in regimes where kinetic barriers and condensation effects become important.
Acknowledgments
The author thanks Prof. H.A. Stone for valuable conversations. This work was supported by the National Science Foundation under award CBET-2417797. The theoretical model development for this work was supported by the Center for Mesoscale Transport Properties, an Energy Frontier Research Center funded by the U.S. Department of Energy, Office of Science, Basic Energy Sciences, under award DE-SC0012673.